NAOJ GW Elog Logbook 3.2

Last entry about OPO alignment is done with the transmission of OPO seperated by normal 50:50 BS. Today I replaced it with a BSY11(dichroic mirror reflects infrared). The first time we align OPO with a temperature of 6.740kOm, but now I found I cannot find green with this tempereture. So I did a scan of temperature, I found green transmission in another temperature. The temperature now is 7.198kOm. (See attached picture 1)This is a roughly fine value, we will find the best value after the lock of OPO. See attached picture 2 of the green I found. And picture 3 of it on the screen. I check the value of infrared transmission again today, which is around 0.4mW. It is similar with last week's condition.

Participants: Yuhang, Eleonora
After achieving the OPO internal alignement (entry #963) we removed the set up we used for this activity (periscope etc..) and we placed the OPO in the final configuration on the bench. (See optical scheme on the wiki updated by Yuhang )
The first rough aligment was done with low power p-pole beam, placing the OPO in order to have its reflection superposed to the incoming beam and then using the last two steering mirrors for the fine alignment.
At the begining we had some trouble with the piezo: since we were only able to get some slowly drifting mode of the camera in trasmission, we suspected that the piezo was not working.
First we check the piezo driver (which was fine), then we check the small box containing electrical connection between BNC from piezo driver and the piezo wires. As soon as we opened the box the piezo start working again. We observed that it was very sensitive to touch. To be checked further.
In the end we could achieve a good aligment, using CC beam (s-pol). See Pic.1. Anyway the trasmission seems much lower than what we expect.
We had only 0.5 mW trasmitted power over 218 mW of the input beam. That is less than 0.25%.
Assuming R1 = 92% and R2 = 99.75%, we should find T = T1*T2 / (1-r1*r2)^2 = 10%
Are the values of R1 and R2 correct? Is the computation of T correct?
This measurement has been done with no temperature control, anyway once we set it on the maximum of the TEM00, transmission seems to be resonably stable to perform the measurement.
Yesterday, I got more information from Matteo. The reflectivity of crystal should be 99.975%. And for in-coupling mirror, it is 92%. By using formula, T=T1*T2/(1-r1*r2)^2, the updated transmission should be 1.192% without considering losses. Now what we found, transmission of 0.25%, seems resonable.


After align OPO with CC beam, I did the alignment with p-pol beam in this morning. As soon as I launch p-pol beam into OPO, I got the attached figure 1.
We know the p-pol power we sent to OPO is only 5.9mW. The OPO choose s-pol to resonant inside the crystal, so we cannot have green if use this p-pol. But I can see signal for infrared on the photodiode. I use this photodiode signal to align the p-pol beam. The steering mirrors I used are the two steering mirrors we put before combinging p-pol and cc. The alignemt result is attached with scan frequency of 60Hz and amplitude of 3Vpp, the signal is shown on the oscilloscope.
After align them, I measured the power I got for transmission, which is 2.5uW. This is very small so I can use camera directly to see it. I put the camera just in front of photodiode. While scanning the PZT with frequency of 50mHz and 3Vpp ramp signal, I can see the modes I have now. Most of them are pitch higher order modes.
This is the best I can do by moving that two steering mirrors. I still don't know why I cannot improve it anymore.

Participants: Matteo, Yuhang, Eleonora
After aligning the crystal along the beam axis, as reported in entry #959, we put the incoupling mirror and align the whole cavity. The aligment has been optimized by maximazing the TEM00 mode from AUX1(Choerent control laser), while the cavity was scanned driving the piezo of the incoupling mirror. The final result is quite good as it can be seen from Fig 1. At the beginning we observed that two different TEM00 mode were resoanting in the same FSR. We found that this was caused by a laser mode hop and was solved by slighly changing the laser temperature.
We also set the crystal temperature control in order to produce green light. The best value was found to be about 6.740 kOhm, but a further optimization will be done once the cavity is locked. The green light produced was transmitted by the dichroic after the cavity and observed with a camera.
We set the PID gains of the temperature controller to the same values of those used for SHG and the control seems to work fine.

Since my result is different from Marc's result, I did calculation again. I found a mistake in my calculation of ABCD matrixs.
From the calculation point of view, Marc's result is correct.
I will check in actual case to see if the calculation aggres with calculation or not. As I have already mentioned in the meeting, we can see the reflected beam is shaking while the filter cavity is locking. So if they don't agree with the actual case, I think the discrepancy comes from the beam shaking.

I measured the profile of the three beams. See the attached plot.
In order to match the waists with the crossing point, I have to move the focusing lens of the 1310nm probe forward by about 25mm, but there is no space.
To make space I'm thinking about adding a 90:10 BS (I have a spare) and change the optical path as shown in the draft design attached.

Yesterday 27 Aug we made another try to move end mirror picomotors. We easily recovered the condition when the beam passing through the hole of the second target was reflected by the end mirror on the back of the target, but displaced few cm in yaw from the center. Unfortunately the yaw picomotor is still stuck and we could not recover the alignement. We are probabably left with no other option that opening the chamber.

Participants: Yuhang, Eleonora, Matteo
The transmission found before seems to come from a beam passing through a gap beside the crystal (it was too bright and the reflection was very small and with a strange shape)
Therefore we have restarted the alignment from the beginning. We moved the the 9071 newport Four-Axis Tilt Aligner under the OPO following the usual procedure:
1. Make the reflection superpose the incoming beam
2. Look for a good transmission, always taking care to keep the reflection superposed to the incoming beam.
After may tries we could find a resonably good shaped transmitted beam with a power of 0.047mW. The input power was 210mW. This means the reflectivity is 99.977%, which is in agreement with what we expect.
We put a dichroic mirror after the OPO. The reflection is focused on a photodiode while the trasmission is sent to a camera. Before installing the OPO we have marked the position of the transmitted beam on the screen in order to use it as a reference for the alignment.
Unfortunaltely the tramission after the dichroic mirror is too low and we cannot see any light on the camera, therefore we were not able to use the reference we took for the transmitted beam.

There were few mistakes made on this entry corrected in this one.
Attached to this entry is the proper fit of beam after the lens (previously a wrong wavelength was used for the plot).
The mean profile was used (w0 = 113.37 um 0.6981 m after the f = 100mm lens used for the characterization [lens is 40 cm after the Faraday Isolator])
The beam parameter is the following : w0 = 18.921 um @ 0.1168 m before the lens ie roughly 0.28320m after the faraday isolator.

I assembled the components for the optical path of the 1310nm probe:
- Fiber collimator. Output mostly p-polarized.
- Polarizing PBS. Transmission: ~90mW, Reflection ~6mW. The reflection goes to a beam dump.
- Half-Wave Plate
- Polarizing PBS rotated so that the reflection goes upward (to a high power beam dump) and the transmission is s-polarized
- Focusing lens f=300mm
- non-polarizing BS 90(R):10(T).
- Transmitted beam (10%):
- Gold coated mirror
- Converging lens f=50mm
- PD DET10N mounted on a XY lens mount
- Reflected beam (90%)
- Gold coated prism mirror.
- To the crossing point...
Then I aligned the beam to impinge at 0.135rad on the sample at the crossing point. The pump is at 2deg (0.035rad) to avoid the etalon effect on the thin samples, this makes the angle between the 1310nm probe and the pump 0.1rad.

- Aligned the imaging unit. Using a blade at about 1cm from the crossing point. I moved the lens on the imaging unit until I saw a sharp image of the blade at the detector position.
- Using the surface reference sample I aligned the pump with the probe to maximize the signal. Then I maximized the signal finely adjusting the position of the whole imaging unit.
- I noticed that there was some noise (around 300uV in the AC signal) with a constant phase. It came from the chopper vibration. So I put some sound absorbing panels and the noise reduced to about 50uV.
- I made a scan of the bulk reference sample with a pump power of 32mW
- I made a scan of the tama-size sapphire and I got the same absorption as before changing the pump size.
- Compared the scans of the bulk reference sample with small pump and large pump. They give the same value (unexpected) but they have a different shape. To be understood.

participaint: Yuhang and Matteo
After the pre-alignment of crystal, I did final algnment yesterday. The method is to move Newport 9071, vertically and herizontally.
The experience is to move two of those four adjustment screws the OPO hole where the beam comes in. The requirement of this adjustment is to make the reflection beam overlap with incident beam. However, at the same time, pay attention that the whole OPO housing should be bascially looking like not tilt. Because we know the beam is bascially aligned pretty well.
During the alignment, I found the reflection looks like the attached figure 1(for the good alignment).
Then we can move the two another side sctews. At the same time, look at the transmission, at a certain point, we can see the trasmission is bright and looks round.
We still need to do aligment of the combination of two horizental screws to make it look better. While doing this, I found the beam is cutted by the crystal after move the incident tilt screw. This means the beam we are looking at is what we want to find. While moving the outgoing tilt screw, we can see the beam is cutted and then many small blocks of lights. This looks like many reflectivity of beam inside the crystal.
In the end, I got the transmission as in attached figure 2 and 3.
Participants: Yuhang, Eleonora, Matteo
The transmission found before seems to come from a beam passing through a gap beside the crystal (it was too bright and the reflection was very small and with a strange shape)
Therefore we have restarted the alignment from the beginning. We moved the the 9071 newport Four-Axis Tilt Aligner under the OPO following the usual procedure:
1. Make the reflection superpose the incoming beam
2. Look for a good transmission, always taking care to keep the reflection superposed to the incoming beam.
After may tries we could find a resonably good shaped transmitted beam with a power of 0.047mW. The input power was 210mW. This means the reflectivity is 99.977%, which is in agreement with what we expect.
We put a dichroic mirror after the OPO. The reflection is focused on a photodiode while the trasmission is sent to a camera. Before installing the OPO we have marked the position of the transmitted beam on the screen in order to use it as a reference for the alignment.
Unfortunaltely the tramission after the dichroic mirror is too low and we cannot see any light on the camera, therefore we were not able to use the reference we took for the transmitted beam.

Participants: Yuhang, Eleonora
We have tried to recover the cavity alignment after TAMA blackout.
1) We have reset all the local controls. For the input and end mirror we have reset the SR560 which filters the error signals (2nd order lowpass filter with cut-off frequency 100 Hz and amplification factor 100)
2) We observed that the reference out of BS chamber, accunting for PR position, was very far from the good position, especially in pitch. Since the local control range was too small, we used picomotors to recover it. Pic 2 shows the reference after moving PR.
3) We had to move also the BS picomotors to recover the good position of the beam on the two in vacuum targets and make the beam to reach the screen after the end mirror, used as a reference for the alignement.
4) At this point we aligned the input mirror in order to have the reflected beam superposed on the input beam.
5) The last step is to align the end mirror. The standard procedure, when no flashes are visible, consists in raising the second target and making the beam to pass through its hole. Than the end mirorr is aligned by making the reflected beam to hit the hole of the target from the back. The reflection can be observed by monitoring the rear side of the target. Unfortunatly the range of the local control was not enough and the end mirror picomotor got stuck in yaw so we were not able to align the end mirror and to see any flash. Pic 2 shows the rear side of the second target, with the reflected beam offset in yaw with respect to the hole of few cm.
SOME DETAIL ON END MIRROR PICOMOTOR: We used both the computer and the joystick to move it. Pitch was always fine. Yaw was moving at the beginning but after a while it got stuck. We tried to switch the driver on and off many times, to swap the channels but it didn't solve the problem.
Since we moved the end picomotor in pitch we observed a higher movement of the mirror with typical peaks in the spectrum which are not normal mode of the suspension (see pic 3). It made me suspect that the intermediate mass is touching the frame of the damping magnets. We already observed something similar in PR suspension (see entry 239 and confront the spectra).
Even if the damping loop is effective in reducing the mirror motion, this should be possibliy solved.
HOW TO GO ON: according to our experience, picomotors which are stuck are likey to start working again after a while. So we will try to move it again in the next days, before considering to open the chamber.
Yesterday 27 Aug we made another try to move end mirror picomotors. We easily recovered the condition when the beam passing through the hole of the second target was reflected by the end mirror on the back of the target, but displaced few cm in yaw from the center. Unfortunately the yaw picomotor is still stuck and we could not recover the alignement. We are probabably left with no other option that opening the chamber.

SOUTH tunnel:
- Dehumidifier: all ON, all green light
- Light: all working
SOUTH end:
- Air conditioning: working
- light: all working
- phone: working (3472)
WEST tunnel:
- Dehumidifier: all ON, 9/10 green light, N.5 (last from the central area before the mid-arm) is ON but red light.
- Light: all working, 3 fluorescent tube were substituted today.
WEST end:
- Air conditioning: NOT working (further inspection needed)
- light: all working
- phone: NOT working


Participaint: Marc, Matteo and Yuhang
Since we have finished the lens installation for OPO, we arrived the alignment of OPO.
Last week, me and Marc did a similar work, including the preparation of periscope, camera, iris, OD filter and the alignment of the beam(make it hight of 140mm). However, the elevation of beam costs too much space. It makes the beam waist locate just after the periscope. After that we decide to alignment the OPO without the translation/tilt stage. However, the work seems impossible without that stage.
Yesterday, we changed the set of stage(make it hight of 111mm). This saves us a lot of space. We also did the beam measurement and find the beam is desirable.(as in the attached figure) We set up the periscope and did the pre-aligment of OPO. The pre-alignment is to make the reflection of OPO roughly overlap with incident beam. The fine-alignment will be done today by using translation/tilt stage. For the components after OPO, we set up a dichroic mirror instead of OD filter to damp the beam and also get the reflection information. For the transmission of dichroic, we use a camera to monitor. We also align it before putting OPO. And made a circle to mark it on the screen.(as in the attached figure). We also set up a photo diode for the reflection of dichroic mirror.
The video of check the reflection is here. https://drive.google.com/file/d/1-ZfhVS0iZDxF7kWpSCrn2QO5ORTTX4xU/view?usp=sharing
The whole set up is attached as well.

Therefore I decided to put a larger breadboard and realign everything.
To do that I maximized the power transmitted by the pinhole in two positions at 2deg from the center (start and end). When the pinhole is at the start position I aligned moving the steering mirror on the optical table. When the pinhole was at the end position I aligned moving the mirror on the periscope (usual procedure to align a beam to make it pass through two points).

When we went to the end room with Takahashi-san, we saw that the south end room (the one we're currently using) was hot and humid.
The power was only off since one stormy day so it can't explain this.
It seems that every now and then the air conditioning system stop in both end rooms.
Seeing the damages in the west end room it could be useful to go from time to time to check if the air conditioning is on in both end rooms.

We slightly moved the positions of the last lens of the OPO telescope as well as the one of the EOM.
We obtain the 2 characterizations attached to this entry.
The nominal value of the beam inside the OPO should be 36 um so we are now within +/- 1.3 um of the nominal value.
The OPO telescope was adjusted so to have the mean beam radius of 36 um.
The x-axis "0" is the last hole of the rail.
The 2 beams are still superposed as we could see on the beam profiler.